I need to solve this with a equation.
A student takes a test where two points are given for every correct answer and 3 points are deducted for every wrong answer. If the student scores 40 points on a 30-question test, how many questions did she answer correctly?

Answer by checkley75(3666)

2X-3(30-X)=40
2X-90+3X=40
5X=130
X=130/5
X=26 CORRECT ANSWERS.
30-26=4 INCORRECT ANSWERS.
PROOF
2*26-3*4=40
52-12=40
40=40